OpenStudy (anonymous):

When rolling a number cube with sides labeled 1 through 6, what is the probability of rolling a number greater than 3? Express your answer in simplest form. Question Resources

OpenStudy (anonymous):

1/2

OpenStudy (welshfella):

requred probability = number of ways to get number > 3 / number of possible outcomes = 3/6 = 1/2

OpenStudy (anonymous):

k